OpenStudy (anonymous):

Using a coupon for 15% off, Joseph paid $1,700 for a new TV. What Was The Regular Price Of The TV?

OpenStudy (shadowfiend):

So what you're saying is that 15% of \(x\) price is $1700. Remember that 15% is the same as 15 / 100. We can set that up like this: \[\frac{15}{100} = \frac{1700}{x}\] Then you can solve for \(x\). You can make it even easier if you flip both sides: \[\frac{100}{15} = \frac{x}{1700}\]

OpenStudy (anonymous):

This is not really the case. I think there is a tiny mistake. What really happens is Joseph gets a discount of 15% on the tv and hence only pays $1.700. So what really happens is J pays only 100%-15%= 85% of the real price. Now if the TV costs $100, J is paying only $85. Now let the real value of the TV be x. We should now see that \[ (100\85)=(x \ 1700) \] Which gives us x=2000 Hence the real value of the tv is $2000. Let me verify. 15% of $2000= 2000* 15/100 = $300 Hence J only pays 2000-300=1700 Hope you understand...

OpenStudy (anonymous):

100 / 85 = x/ 1700 Small typo! Not x1700!

OpenStudy (shadowfiend):

Oh dear, you're totally right samith! Sorry!